# Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance in reasoning tasks but face limitations in mathematical and complex logical reasoning. Existing methods to improve LLMs’ logical capabilities either involve traceable or verifiable logical sequences that generate more reliable responses by constructing logical structures yet increase computational costs, or introduces rigid logic template rules, reducing flexibility. In this paper, we propose Reversal of Thought (RoT), a novel framework aimed at enhancing the logical reasoning abilities of LLMs. RoT utilizes a Preference-Guided Reverse Reasoning warm-up strategy, which integrates logical symbols for pseudocode planning through meta-cognitive mechanisms and pairwise preference self-evaluation to generate task-specific prompts solely through demonstrations, aligning with LLMs’ cognitive preferences shaped by Reinforcement Learning with Human Feedback (RLHF). Through reverse reasoning, we ultilize a Cognitive Preference Manager to assess knowledge boundaries and further expand LLMs’ reasoning capabilities by aggregating solution logic for known tasks and stylistic templates for unknown tasks. Experiments across various tasks demonstrate that RoT surpasses existing baselines in both reasoning accuracy and efficiency.
@@ 1 Introduction
Recent studies have advanced CoT to guide LLMs, mainly through either multi-step prompting such as introducing planning-and-solve Plaat et al. (2024); Yang et al. (2024a), self-consistency Narang et al. and recursive reasoning process Lee and Kim (2023); Yu et al. (2024) through Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024), or multi-role Zhang et al. ; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased the computational cost of inference or API access cost due to the multi-step inference. Buffer-of-Thought (BoT) Yang et al. (2024a) attempts to reduce thinking steps by leveraging Retrieval-Augmented Generation (RAG) to retrieve gold thought templates from the buffer. However, it sacrifices flexibility due to the initialization of pre-set manual thought templates. Therefore, achieving accurate reasoning in LLMs while minimizing resource consumption remains a significant challenge.
@@
To address above limitations, we propose Reversal of Thought (RoT), a novel framework that enables LLMs to explore cognitive preference on logical pseudocode solely using reverse prompting with given demos without additional task-related affirmations, as depicted in Figure 1. Our key contributions are as follows:
  • To the best of our knowledge, we are the first to introduce a reversal reasoning for cognitive preference that enhances logical reasoning in LLMs by combining meta-cognitive with cognitive preference, resulting in a more cost-efficient and error-resilient framework for complex tasks.
@@
  • We introduce logical symbols in pseudocode for algorithmic planning and problem-solving, guided by meta-cognitive mechanisms, which improves the LLMs’ capacity for structured and accurate logical reasoning.
@@ 2.1 Chain-of-Thought (CoT) Prompting
Chain-of-Thought (CoT) prompting Wei et al. (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic. Recent studies primarily aimed at improving logical accuracy by introducing more external validation such as self-consistency Narang et al. ; Yu et al. (2024) or more hierarchical information such as Least-to-Most Zhou et al. , Cumulative-Reasoning Zhang et al. and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors or poor flexibility. Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024) and Buffer-of-Thought (BoT) Yang et al. (2024a). However, an efficient logical reasoning method that strikes a balance among logical accuracy, flexibility, and cost has yet to be discovered. Our method is activated through meta cognition Fleur et al. (2021) by introducing reverse reasoning to form effective LLMs-taste prompts within cognitive preference Uddin (2021) for plan-and-solve with logical pseudocode at least.
@@ 2.2 Knowledge Boundary for Enhancing Large Language Models
Integrating knowledge boundary within LLMs has emerged as a prospective strategy for enhancing their ability to avoid reasoning hallucinations of unknown knowledge through knowledge boundary constraints which requires additional algorithmic efforts Yin et al. (2024), external graph knowledge Tian et al. (2024), and training consumption Sun et al. (2024). Additionally, they focus on avoiding responses to unknown or incorrect prompts rather than proposing bold and proactive solutions to expand knowledge boundary in a heuristics without training.
@@ 3.1 Overview
As the aforementioned wisdom related to human cognitive learning implies, merely telling or teaching is inadequate Bao et al. (2024). Moreover, most LLMs have undergone extensive pre-training Achiam et al. (2023); Bai et al. (2023); Dubey et al. (2024) and reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), instilling in LLMs a propensity for specific cognitive patterns, which manifests in two progressive layers of LLMs-taste description: (1) Stylistic template: encompassing grammatical and syntactic structures in descriptions. (2) Solution logic: comprising problem-solving reasoning and methodological cues. Therefore, Reversal of Thought (RoT) involves answering the following two research questions (RQs):
@@ 3.2 Reverse Reasoning with Meta-cognition
Inspired by RLHF Ouyang et al. (2022) utilizing preference data, and to derive high-cognitive preference prompt $P^{*}$ that enhance logical reasoning in LLMs, we propose a Preference-Guided Reverse Reasoning (PGRR) framework (detailed in alogrithm 1) mapping input-output demonstrations $D$ from an initial prompt $P$ to an optimal LLM-taste prompt $P_{opt}$.
@@
We query the LLM $\mathcal{M}_{\text{LLM}}$ with a reversal prompt and demonstrations $\{P_{r},D\}$ (detailed in figure 3) for $warm$ iterations, generating a set of candidate cognitive preference responses $R=\{R_{1},R_{2},\dots,R_{warm}\}$ and their corresponding average probabilities $P_{res}^{i}$:
@@
where $\mathcal{M_{LLM}}(Prompt_{eval},R_{i+1},R_{i})$ represents that require LLM to select more preferred data through $P_{eval}$ with a structure as "Please choose your more preferred instruction (A/B): (A).$R_{i+1}$; (B).$R_{i}$".
Following the principle of preference transitivity Liu et al. (2024), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
@@
Within reverse reasoning, we further follow meta-cognitive Suzgun and Kalai (2024) using plan-and-solve by integrating logical algorithm pseudo-code to improve reasoning comprehension. And we incorporate fundamental mathematical logic symbols, including logical operators, quantifiers, inequalities and conditional statements, to facilitate model reasoning detailed in Figure 3.

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable performance in reasoning tasks but face limitations in mathematical and complex logical reasoning. Existing methods to improve LLMs’ logical capabilities either involve traceable or verifiable logical sequences that generate more reliable responses by constructing logical structures yet increase computational costs, or introduces rigid logic template rules, reducing flexibility. In this paper, we propose Reversal of Thought (RoT), a plug-and-play and cost-effective reasoning framework designed to enhance the logical reasoning abilities of LLMs during the warm-up phase prior to batch inference. RoT utilizes a Preference-Guided Reverse Reasoning warm-up strategy, which integrates logical symbols for pseudocode planning through meta-cognitive mechanisms and pairwise preference self-evaluation to generate task-specific prompts solely through demonstrations, aligning with LLMs’ cognitive preferences shaped by RLHF. Through reverse reasoning, we utilize a Cognitive Preference Manager to assess knowledge boundaries and further expand LLMs’ reasoning capabilities by aggregating solution logic for known tasks and stylistic templates for unknown tasks. Experiments across various tasks demonstrate that RoT surpasses existing baselines in both reasoning accuracy and efficiency.
[…]
[ roundcorner=5pt, backgroundcolor=UserExampleBg, linecolor=UserExampleTitle, outerlinewidth=0.5pt, frametitlebackgroundcolor=UserExampleTitle, frametitlefont=, ]user_example
@@ 1 Introduction
Recent studies have advanced CoT to guide LLMs, mainly through either multi-step prompting such as introducing planning-and-solve Plaat et al. (2024); Yang et al. (2024a), self-consistency Narang et al. ; Wang et al. (2024) and recursive reasoning process Lee and Kim (2023); Yu et al. (2024) through Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024), or multi-role Zhang et al. ; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased inference cost due to the multi-step inference. Buffer-of-Thought (BoT) Yang et al. (2024a) attempts to reduce thinking steps by leveraging Retrieval-Augmented Generation (RAG) to retrieve gold thought templates from the buffer. However, it sacrifices flexibility due to the initialization of pre-set manual thought templates. Therefore, achieving accurate reasoning in LLMs while minimizing resource consumption remains a significant challenge.
@@
To address above limitations, inspired by meta-cognition Fleur et al. (2021) and and cognitive preference Uddin (2021); Zhou et al. (2023); Margatina et al. (2023), we propose Reversal of Thought (RoT), a plug-and-play and cost-effective framework that enables LLMs to explore cognitive preference on logical pseudocode solely using reverse prompting with given demos without additional task-related affirmations, as depicted in Figure 1. Our key contributions are as follows:
  • To the best of our knowledge, we are the first to introduce a reversal reasoning for cognitive preference that enhances logical reasoning in LLMs by combining meta-cognitive with cognitive preference, resulting in a more modular and cost-efficient framework for complex tasks.
@@ 2.1 Chain-of-Thought (CoT) Prompting
Chain-of-Thought (CoT) prompting Wei et al. (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic. Recent studies primarily aimed at improving logical accuracy by external validation mechanisms like symbolic reasoning Cai et al. (2023); Pan et al. , stepwise verification including self-consistency Narang et al. ; Yu et al. (2024); Wang et al. (2024), self-refine Madaan et al. (2024), self-reflection Renze and Guven (2024) and more hierarchical information such as Least-to-Most Zhou et al. , Cumulative-Reasoning Zhang et al. and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors Bao et al. (2024) or poor flexibility Yang et al. (2024a). Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024) and Buffer-of-Thought (BoT) Yang et al. (2024a). However, an efficient logical reasoning method that strikes a balance among reasoning accuracy, flexibility, and cost has yet to be discovered. Our method is activated through meta cognition Fleur et al. (2021) by introducing reverse reasoning to form effective LLMs-taste prompts within cognitive preference Uddin (2021) for plan-and-solve with logical pseudocode at least.
@@ 2.2 Knowledge Boundary for Enhancing Large Language Models
Integrating knowledge boundary within LLMs has emerged as a prospective strategy for enhancing their ability to avoid reasoning hallucinations of unknown knowledge through knowledge boundary constraints which requires additional algorithmic efforts Yin et al. (2024); Chen et al. , external graph knowledge Tian et al. (2024), and training consumption Sun et al. (2024). Additionally, they focus on avoiding responses to unknown or incorrect prompts rather than proposing bold and proactive solutions to expand knowledge boundary in a heuristics without training.
@@ 3.1 Overview
As the aforementioned wisdom related to human cognitive learning implies, merely telling or teaching is inadequate Bao et al. (2024). Moreover, most LLMs have undergone extensive pre-training Achiam et al. (2023); Bai et al. (2023); Dubey et al. (2024) and reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), instilling in LLMs a propensity for specific cognitive patterns, which manifests in two progressive layers of LLMs-taste description: (1) Stylistic template: encompassing grammatical and syntactic structures in descriptions for thinking problems. (2) Solution logic: comprising problem-solving reasoning and methodological cues. Therefore, Reversal of Thought (RoT) involves answering the following two research questions (RQs):
@@ 3.2 Reverse Reasoning with Meta-cognition
Inspired by RLHF Ouyang et al. (2022) utilizing preference data, and to derive high-cognitive preference prompt $P^{*}$ that enhance logical reasoning in LLMs, we propose a Preference-Guided Reverse Reasoning (PGRR) framework (detailed in alogrithm 1) mapping input-output demonstrations $D$ from an initial prompt $P$ to an optimal LLM-taste prompt $P_{opt}$ from prompt candidates.
@@
We query the LLM $\mathcal{M}_{\text{LLM}}$ with a reversal prompt and demonstrations $\{P_{r},D\}$ (detailed in figure 4) for $warm$ iterations, generating a set of prompt candidates for solution $R=\{R_{1},R_{2},\dots,R_{warm}\}$ and their corresponding average probabilities $P_{res}^{i}$:
@@
where $\mathcal{M_{LLM}}(P_{eval},R_{i+1},R_{i})$ represents that require LLM to select more preferred data through $P_{eval}$ with a structure as "Please choose your more preferred instruction (A/B): (A).$R_{i+1}$; (B).$R_{i}$".
Following the principle of preference transitivity Liu et al. (2024b), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
@@
Within reverse reasoning, we further follow meta-cognitive Suzgun and Kalai (2024) using plan-and-solve by integrating logical algorithm pseudo-code to improve reasoning comprehension. And we incorporate fundamental mathematical logic symbols, including logical operators, quantifiers, inequalities and conditional statements, to facilitate model reasoning detailed in Appendix A.2.
3.3 Cognitive Preference Manager
Algorithm 2 Cognitive Preference Manager (CPM)
1:  $P_{task}\leftarrow P,P^{*}_{task}\leftarrow P^{*}$
3:  if $s\geq\delta$ then
5:     Aggregate relevant task-specific knowledge.
6:     return  $P_{final}$.
7:  else
8:     UnKnown detected: Adapt and expand $P_{task}$.
10:     return  $P_{final}$.
11:  end if
Cognitive Preference Manager.
By efficiently evaluating cognitive results, CPM integrates alternative aggregation strategies for $\mathcal{M_{LLM}}$ based on $S_{cog}$ 111we use $S_{cog}$ to conduct reverse evaluation of GPT-4 in our experiments., as detailed in appendix A.3: (1) Solution logic aggregation for known tasks: $\mathcal{M_{LLM}}$ merges beneficial aspects from the original prompt $P$ with the LLM-taste prompt $P^{*}$ to create the final prompt $P_{\textit{final}}$. (2) Stylistic template aggregation for unknown $\mathcal{M_{LLM}}$ extracts a cognitive preference template for thinking $T$ from the incorrect context in the LLM-taste prompt, and integrates meta-cognitive elements from the original prompt $P$ into $T$ to construct the final prompt $P_{final}$.
$answer=\mathcal{M_{LLM}}(P_{final},problem)$ (8)
4 Experiments
4.1 Datasets and Tasks
Following Yao et al. (2024); Suzgun and Kalai (2024); Yang et al. (2024a), we evaluate RoT across eight tasks (detailed in Appendix A.1): (1) Mathematical Reasoning:Game of 24 Yao et al. (2024); Xiang et al. (2025), Multilingual Grade School Math Shi et al. (2022); (2) Domain Knowledge: Python Puzzles Schuster et al. (2021), Geometric Shapes, Multi-Step Arithmetic Two, Word Sorting, and Checkmate-in-One Srivastava et al. (2023); (3) Creativity: Shakespearean Sonnet Writing Suzgun and Kalai (2024).
4.2 Baselines
4.3 Experiment Setup
To ensure fair comparisons with previous methods, following Liu et al. (2024a); Yao et al. (2024); Zhang et al. ; Suzgun and Kalai (2024), we utilize two LLMs: GPT-3.5-turbo and GPT-4, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study. For the $warmup$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API. Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency. For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^2] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al. (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
4.4 Evaluation Metrics
Baselines Game of 24 Geometric Shapes Multi-Step Arithmetic Word Sorting Checkmate-in-One Python Puzzles MGSM Sonnet Writing
GPT-3.5-turbo
CoT (NeurIPS) Wei et al. (2022) 1-shot 26.3 50.2 73.9 71.9 25.7 30.1 75.0 58.6
2-shot 29.6 53.2 76.2 74.6 25.4 30.6 77.4 61.3
Meta-prompting Suzgun and Kalai (2024) 1-shot 59.6 69.5 83.1 90.0 47.4 40.2 80.3 69.5
2-shot 60.2 71.0 85.1 91.5 48.0 41.1 81.5 71.4
ToT (NeurIPS) Yao et al. (2024) 1-shot 66.7 56.2 81.9 89.2 44.2 38.2 78.3 63.5
2-shot 67.3 58.5 82.3 90.3 45.2 39.5 79.4 65.3
GoT (AAAI) Besta et al. (2024) 1-shot 69.7 58.6 79.1 88.8 45.4 35.3 80.3 62.2
2-shot 70.4 58.9 79.0 89.5 46.0 36.7 81.0 63.7
BoT (NeurIPS) Yang et al. (2024a) 1-shot 74.7 83.1 87.6 93.2 67.9 48.2 81.1 71.1
2-shot 75.5 84.3 88.3 94.4 68.5 48.0 82.3 73.0
RoT (Ours) 1-shot 82.8 88.4 89.2 95.2 71.5 50.2 82.3 75.5
2-shot 87.8 88.7 89.5 95.6 72.6 50.4 84.7 75.8
GPT-4
CoT (NeurIPS) Wei et al. (2022) 1-shot 31.3 57.4 83.1 80.7 35.8 35.3 83.9 66.3
2-shot 32.7 60.9 85.5 82.7 36.7 35.9 84.3 70.2
Meta-prompting Suzgun and Kalai (2024) 1-shot 64.6 76.7 89.6 97.6 58.2 45.0 85.1 78.7
2-shot 66.3 78.2 90.7 98.4 58.9 45.6 85.5 79.8
ToT (NeurIPS) Yao et al. (2024) 1-shot 73.7 58.6 88.8 95.2 49.8 43.8 84.7 68.7
2-shot 74.5 60.1 90.3 96.0 48.8 45.2 86.3 69.8
GoT (AAAI) Besta et al. (2024) 1-shot 74.7 56.2 87.6 95.6 49.0 42.6 85.9 68.3
2-shot 73.5 57.7 88.7 96.8 51.2 43.5 86.7 62.5
BoT (NeurIPS) Yang et al. (2024a) 1-shot 82.8 90.4 94.8 99.2 87.1 51.8 87.9 79.1
2-shot 83.7 90.7 96.8 99.6 88.3 52.8 87.5 79.8
RoT (Ours) 1-shot 97.0 94.8 98.4 99.6 91.2 54.6 88.7 89.2
2-shot 98.0 95.2 99.2 100.0 92.0 57.3 90.0 92.0
Knowledge Boundary and Cognitive Preference Consistency.
As described in Section 3.3, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$. We first evaluate GPT-3.5-turbo & GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^3] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot. and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
Reasoning Accuracy.
Following Suzgun and Kalai (2024); Yao et al. (2024), we introduce a LLM444we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong). We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
Reasoning Efficiency.
Following Yang et al. (2024a), we evaluate reasoning efficiency in terms of complexity by calculating $T$, calculated as the average time spent per task across all samples in the test dataset:
$T=\frac{1}{N}\sum_{i=1}^{M}\sum_{j=1}^{N}T_{ij}$ (9)
5 Results and Discussion
5.1 Knowledge Boundary and Cognitive Preference Consistency
Task 1-shot 2-shot
Kno. Unkno. Kno. Unkno.
Game of 24
MGSM (avg)
Multi-Step Arithmetic
WordSorting
Python Puzzles
Geometric Shapes
Checkmate-in-One
Sonnet Writing
Table 2: Knowledge Boundary under Reversal Reasoning for GPT-3.5-turbo and GPT-4 with identical results. Kno. for Known, Unkno. for Unknown.
Knowledge Boundary.
As shown in Table 2, our findings provide preliminary support for our hypothesis concerning cognitive knowledge boundaries of LLMs in reverse reasoning across current task benchmarks in one-shot and two-shot settings Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024). Notably, GPT-3.5-turbo & GPT-4 excel in structured reasoning tasks such as Game of 24, Geometric Shapes, and Checkmate-in-One, consistently achieving strong results in both one-shot and two-shot settings. In contrast, tasks such as MGSM (avg) and Python Puzzles fall into an unknown category in the one-shot scenario, stemming from multi-source problems that are challenging to make reverse reasoning without sufficient context, indirectly supporting the rationale for our CPM approach.
Cognitive Preference Consistency.
Our human evaluation of $\textit{Con}_{cog}$ yielded high scores of 4.32 in the 2-shot setting and 4.01 in the 1-shot setting, further validating that LLMs exhibit cognitive preferences shaped by RLHF Ouyang et al. (2022).
5.2 Reasoning Accuracy and Efficiency
(1) RoT can activate LLMs’ reasoning accuracy.
As shown in Table 1, RoT consistently outperforms all baselines across various reasoning tasks, with particularly notable improvements observed in GPT-4. Specifically, compared with the best BoT, RoT achieves significant gains in tasks such as Game of 24 (+17.15% in 1-shot and +17.08% in 2-shot), Geometric Shapes (+4.87% in 1-shot and +4.96% in 2-shot), and Checkmate-in-One (+4.71% in 1-shot and +4.19% in 2-shot), demonstrating our flexibility of leveraging cognitive preference in LLM to activate logic capabilities through reverse reasoning (case study detailed in appendix B). For GPT-3.5-turbo, RoT also demonstrates substantial improvements, such as Game of 24 (+10.84% in 1-shot and +16.29% in 2-shot) and Geometric Shapes (+6.38% in 1-shot and +5.22% in 2-shot), further emphasizing its versatility in activating logic through cognitive preference management.
Figure 3: Inference time comparison
Ablation study Game of 24 Geometric Shapes Multi-Step Arithmetic Word Sorting Checkmate-in-One Python Puzzles×$\times$ MGSM×$\times$ Sonnet Writing
GPT-3.5-turbo
RoT 1-shot 82.8 88.4 89.2 95.2 71.5 50.2 82.3 75.5
2-shot 87.8 88.7 89.5 95.6 72.6 50.4 84.7 75.8
w/o PGRR 1-shot 74.7 72.3 74.3 74.7 65.5 44.2 78.7 69.6
2-shot 75.5 72.6 75.4 76.2 66.9 44.4 79.0 69.8
w/o Logic 1-shot 76.7 80.7 81.5 84.3 66.3 45.8 77.5 72.9
2-shot 76.5 80.2 81.9 85.1 66.5 46.0 78.2 73.0
w/o CPM 1-shot 80.8 86.0 88.4 93.2 70.3 39.8 72.7 74.2
2-shot 81.6 86.7 88.3 93.5 71.0 47.2 80.2 75.0
GPT-4
RoT 1-shot 97.0 94.8 98.4 99.6 91.2 54.6 88.7 89.2
2-shot 98.0 95.2 99.2 100.0 92.0 57.3 90.0 92.0
w/o PGRR 1-shot 77.8 74.7 78.7 79.5 80.7 45.0 85.9 72.7
2-shot 78.6 75.0 79.0 81.0 82.3 46.8 86.0 73.0
w/o Logic 1-shot 85.9 83.1 87.1 89.2 81.5 49.8 85.5 86.1
2-shot 86.7 83.5 87.9 89.5 83.5 50.0 86.7 77.0
w/o CPM 1-shot 92.0 90.0 94.8 97.6 89.2 40.6 84.7 88.8
2-shot 93.9 90.3 95.2 98.4 89.1 51.6 86.3 91.5
(2) RoT demonstrates better tradeoff between reasoning accuracy and efficiency across baselines.
As shown in Figure 3, RoT achieves competitive performance in reasoning efficiency, outperforming baselines and being second only to BoT, while BoT’s dependence on numerous pre-defined golden thought templates limits its flexibility. In contrast, our RoT, as a plugin strategy, emphasizes exploring optimal prompt for solution and task data instantiation after reverse reasoning warm-up.
5.3 Ablation Study
Impact of PGRR.
Excluding Preference-Guided Reverse Reasoning (PGRR) results in a significant reduction in overall task performance, as seen in tasks like the Game of 24 (98.0% to 78.6% for GPT-4, 87.8% to 75.5% for GPT-3.5-turbo) and WordSorting (100% to 81% for GPT-4, 95.6% to 76.2% for GPT-3.5-turbo), indicating w/o PGRR weaken model’s cognition to task-specific requirements.
Impact of Logic.
Removing mathematical logic from RoT leads to notable declines in tasks requiring structured problem-solving, such as Multi-Step Arithmetic (99.2% to 87.9% for GPT-4, 89.5% to 81.9% for GPT-3.5-turbo) and Checkmate-in-One (92.0% to 83.5% for GPT-4, 72.6% to 66.5% for GPT-3.5-turbo), underscoring w/o Logic negatively affects structured problem-solving for complex reasoning.
Impact of CPM.
Lacking Cognitive Preference Manager (CPM) has a particularly pronounced effect on unknown tasks, with Python Puzzles dropping from 54.6% to 45.0% for GPT-4 and 50.2% to 39.8% for GPT-3.5-turbo, while MGSM decreased from 84.7% to 80.2% for GPT-3.5-turbo and from 90.0% to 86.3% for GPT-4, indicating w/o CPM weaken RoT’s flexibility for tackling known tasks and unknown tasks  [^5] Please refer to Appendix B.1.1 for case study on known task and Appendix B.2.1 for case study on unknwon task..
6 Conclusion
In this paper, we propose Reversal of Thought (RoT), a novel and plug-and-play framework to enhance the logical reasoning capabilities of LLMs. By integrating reverse reasoning with meta-cognitive mechanisms and cognitive preference management, RoT improves reasoning accuracy and efficiency while minimizing computational costs, which leverages Preference-Guided Reverse Reasoning and Cognitive Preference Manager, which optimally aligns LLM reasoning processes with their cognitive preferences shaped by their pretraining and RLHF. Comprehensive experiments across diverse reasoning tasks demonstrate that RoT consistently outperforms state-of-the-art baselines in both known and unknown task scenarios, demonstrating the potential to expand knowledge boundaries through cognitive preference template. Our research provides valuable insights into future studies focused on further enhancing LLMs’ reasoning capacities by dynamically exploring cognitive preferences for complex reasoning tasks.
Limitations
In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks. Furthermore, we believe that utilizing Reversal of Thought in teacher-student model distillation, which could further amplify the practical value of our approach.
Ethical Considerations
Since the datasets used in our experiments focus on pure mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity, which are all sourced from publicly available datasets Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024); Xiang et al. (2025) and devoid of any personal privacy or sensitive ethical information. Therefore, we do not identify any immediate ethical concerns regarding our current work. Additionally, we conduct human evaluations of cognitive preference consistency with three anonymous professional annotators following our instructions (detailed in appendix C).
Appendix A Experimental Details
A.1 Datasets and Tasks
To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al. (2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al. (2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al. (2024); Xiang et al. (2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24666https://www.4nums.com/game/ . (2) BIG-Bench Suzgun et al. (2023); Srivastava et al. (2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^7] https://huggingface.co/datasets/google/bigbench Srivastava et al. (2023); (3) Python Puzzles Schuster et al. (2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels [^8] https://github.com/microsoft/PythonProgrammingPuzzles;(4) Multilingual Grade School Math (MGSM) Shi et al. (2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al. (2021), featuring translations of a subset of examples into ten diverse languages [^9] https://github.com/google-research/url-nlp/tree/main/mgsm .(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^10] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^11] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
A.2 Prompt for Reverse Reasoning